# CHG Shapley: Efficient Data Valuation and Selection towards Trustworthy Machine Learning

### v2

@@ 3.3 CHG Shapley
Proof see Appendix B
We implemented our CHG Shapley data valuation method (see Algorithm 1) based on the code foundation of OpenDataVal (Jiang et al., 2023) and conducted experiments on the CIFAR-10 embedding dataset on a NVIDIA A40 server. The model used is pre-trained ResNet50 + logistic regression, and we report the results in Table 1.
For noisy label detection tasks and comparisons with other data evaluation methods, see Picture 2. For point removal tasks, see Picture 1.
The experimental results demonstrate that CHG Shapley data valuation method not only exhibits excellent data valuation and noise data detection capabilities, but also achieves remarkable operational efficiency. This is made possible by our research into the utility function that measures the impact of data subsets on model accuracy. In the next section, we will further show its data selection ability based on data valuation and noise label detection.
@@
0:  Training dataset ($N$), model architecture ($\theta_{0}$)
0:  Value of every training data
1:  Initialization: Model parameter and training hyperparameters initialization
2:  for $k$-th training epoch do
3:     Acquire loss and last-layer gradients of each data, i.e., $l_{i},\nabla f(i;\theta_{k})$
4:     Calculate Shapley value of every data point based on Equation 3, with $\alpha=\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)$ and $x_{i}=l_{i}\nabla f(i;\theta_{k})$
5:     Update model parameters $\theta_{k}$ to $\theta_{k+1}$
6:  end for
7:  Return each data’s average Shapley value
@@ 4 Data Selection
In this section, we introduce various implementation strategies and practical techniques to enhance the scalability and efficiency of CHG Shapley. The experiments are conducted on a NVIDIA A40 server.
Last-Layer Gradients: Modern deep learning models possess a vast number of parameters, resulting in high-dimensional gradients. To mitigate this issue, we adopt a last-layer gradient approximation as in(Killamsetty et al., 2021a, b). This technique focuses solely on the gradients of the last layer for neural networks in CHG Shapley, significantly improving the speed of the algorithm and other baseline methods.
Per-Class Approximations: To enhance model performance, we adopt a per-class approach. This method involves caluclate the CHG Shapley separately for each class, which means considering only the data instances belonging to each class in each iteration. This approach not only reduces memory requirements but also decreases computational overhead, thereby improving the overall efficiency of the CHG Shapley-based data selection algorithm.
CHG Shapley-based Data Selection Algorithm: Here, we detail the CHG Shapley Data Selection Algorithm, integrating the aforementioned techniques to ensure efficient and scalable data valuation and selection.
@@
0:  Training dataset ($N$), model architecture ($\theta_{0}$), selection interval $R=20$, the fraction (denoted as $a$) of data available for each epoch.
0:  Model parameters
1:  Initialization: Model parameter and training hyperparameters initialization, and let the training subset $S=\emptyset$.
2:  for $k$-th training epoch do
@@
6:           Acquire loss and last-layer gradients of each data in this class, i.e., $l_{i},\nabla f(i;\theta_{k})$
7:           Calculate Shapley value of every data point in this class (denoted as $N_{c}$) based on Equation 3, with $\alpha=\frac{1}{N_{c}}\sum_{i\in N_{c}}l_{i}\nabla f(i;\theta)$ and $x_{i}=l_{i}\nabla f(i;\theta_{k})$
8:           Add the top $aN_{c}$ data points with the highest Shapley values to the training subset $S$
@@
10:        Calculate the Shapley values of each data point in $S$ with $\alpha=\frac{1}{S}\sum_{i\in S}l_{i}\nabla f(i;\theta)$ and $x_{i}=l_{i}\nabla f(i;\theta_{k})$, and apply minmax normalization to Shapley values to get the weights (denoted as $\{w_{S}\}$) of each data point in $S$
11:     end if
12:     Train model on $S$ using weighted learning with $\{w_{S}\}$, updating its parameters to $\theta_{k+1}$
13:  end for
Results: We report the accuracy and time consumption comparisons in Table 2 under the data pruning setting. CHG Shapley demonstrates particularly strong performance when the selection ratio is small, highlighting its ability to identify high-value data effectively. Additionally, the time consumption remains acceptable even when considering data selection processes, underscoring the high efficiency of CHG Shapley as a data valuation method. This efficiency, combined with its accuracy, verifies that CHG Shapley is a practical and powerful tool for data selection and valuation in large-scale machine learning tasks.
The results presented in Table 3, where 30% of the class labels were randomly replaced with other labels, provide additional evidence of CHG Shapley’s capability to detect mislabeled data. CHG Shapley demonstrates superior performance compared to other methods in this scenario, reinforcing its effectiveness in identifying and addressing label noise.
@@
Table 2: The accuracy (%) and time (s) comparison. All methods are trained with ResNet-18.
@@
Dataset CIFAR10 CIFAR100 Tiny ImageNet Fraction 0.05 0.1 0.3 0.5 0.7 0.05 0.1 0.3 0.5 0.7 0.05 0.1 0.3 0.5 0.7 Accuracy (%) Full 95.51 77.56 62.43 Random 65.90 77.36 89.93 93.11 93.87 22.58 36.51 62.27 69.70 73.04 17.13 27.70 45.17 52.37 57.89 AdaptiveRandom 83.26 88.78 93.88 94.77 94.66 48.52 60.58 72.45 74.97 75.70 37.76 48.50 58.49 60.30 62.18 Glister 85.44 89.92 91.26 91.93 93.04 48.25 64.63 72.72 74.00 75.14 37.17 48.29 56.77 59.13 61.50 GradMatch (PB) 85.17 90.17 93.93 94.42 94.83 50.42 63.53 73.79 75.78 76.20 38.88 49.97 59.79 61.56 62.19 CHG Shapley 87.15 90.88 93.65 94.51 94.58 53.97 64.46 72.54 74.38 75.71 38.92 49.00 58.44 60.68 61.80 Hardness Shapley 54.96 68.21 84.18 87.02 - 34.54 53.11 66.20 69.10 - - - - - - Gradient Shapley 84.87 89.95 93.54 94.08 - 51.15 59.73 70.69 73.88 - - - - - - Time (s) Full 7745 8082 49747 Random 776 1140 1906 6489 8708 805 1645 3839 3170 5859 2742 5184 15323 24475 33910 AdaptiveRandom 502 1522 4037 6815 6574 746 1358 3324 4786 7166 3451 5852 21366 27659 37583 Glister 1600 2736 4713 11853 15740 1247 1885 4778 11762 14611 4382 7693 19210 28289 36558 GradMatch (PB) 1308 2129 4897 7661 9916 804 2192 4077 6095 9185 4455 6904 24238 30814 51780 CHG Shapley 1383 1768 3746 7356 9526 1258 2037 4706 7012 9657 7077 9887 25682 30762 38177 Hardness Shapley 1191 1741 3606 4878 - 1233 1836 3828 5824 - - - - - - Gradient Shapley 1365 1933 2714 4017 - 1427 2025 2730 4032 - - - - - -
Table 3: Comparison of performance under noise label setting.
Dataset CIFAR10 CIFAR100 Fraction 0.05 0.1 0.3 0.5 0.7 0.05 0.1 0.3 0.5 0.7 AdaptiveRandom 57.18 63.21 71.84 75.38 77.41 24.12 32.69 45.36 52.18 53.00 Glister 61.17 70.39 73.73 75.64 76.59 18.64 29.22 45.80 49.78 52.45 GradMatch (PB) 58.38 69.67 72.59 75.41 76.90 25.90 33.84 47.71 52.35 54.29 CHG Shapley 67.43 77.32 85.33 88.83 88.90 35.75 46.41 59.59 61.92 63.36
@@ 5 Discussion and Future Work
In this paper, we define the Conduct of Hardness and Gradient score (CHG score) of a data subset $S$ as $\|\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)\|_{2}^{2}-\|\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)-\frac{1}{S}\sum_{i\in S}l_{i}\nabla f(i;\theta)\|_{2}^{2}$, and use this score based on Euclidean distance to measure the influence of data subsets on model accuracy. However, it’s worth noting that many studies on data selection based on gradients utilize different distance measures between $\frac{1}{N}\sum_{i\in N}l_{i}\nabla f(i;\theta)$ and $\frac{1}{S}\sum_{i\in S}l_{i}\nabla f(i;\theta)$, such as cosine similarity (Zhao et al., 2021), inner product, L1 norm, or their combinations, and then optimize this metric to select data. However, whether these metrics can be used for data valuation remains unknown because, unlike the theoretical support for the aforementioned Euclidean distance, these metrics lack theoretical grounding. Whether cosine similarity, inner product, L1 norm, or their combinations can achieve this goal remains to be investigated.
On the other hand, if cosine similarity and L1 norm are used in our framework, it becomes challenging to derive the corresponding Shapley values. In other words, if the Euclidean distance cannot effectively measure the utility of dataset subsets for model accuracy, the entire data valuation approach proposed in this study becomes computationally infeasible. The root of this problem lies in the complexity of computing Shapley values, which is generally challenging and only tractable when the utility function follows a specific form (such as linear models, Euclidean distance, or inner product).

### v3

@@ 3.3 CHG Shapley-based Data Valuation Algorithm
Proof see Appendix B.
The use of CHG Shapley to calculate the data value of each training data point is outlined in Algorithm 1, and the experimental results for different data valuation methods are presented in Table 1. Detailed experimental results comparing these data valuation methods can be found in Appendix A.
Differences from Data Shapley: Although our approach shares the same underlying principle as Data Shapley, we do not try to approximate it. Data Shapley utilizes a model trained on a subset of the training set, evaluated on an additional validation set, as its utility function. In CHG Shapley, our utility function captures the extent of training loss reduction when using a subset of the training dataset. Furthermore, in Data Shapley, the required size of the validation set can grow by orders of magnitude to achieve the desired utility precision (i.e., the model’s accuracy on the validation set). For example, in a classification task, achieving a precision of 0.001 for the utility function may require 1,000 validation samples, while a precision of 0.0001 could demand 10,000 samples. Another challenge lies in the noise affecting the utility function, which arises from the model training process—particularly due to stochastic gradient descent—as discussed in Data Banzhaf (Wang & Jia, 2023). In contrast, CHG Shapley relies solely on the training set and incorporates both loss and gradient information from the model during training. This richer and more precise information allows CHG Shapley to achieve lower computational costs and higher precision. Moreover, by averaging the Shapley values over multiple epochs, CHG Shapley is somewhat resistant to the impacts of utility noise.
@@
0:  Training dataset $N$, initial model parameters $\theta_{0}$, total epochs $K$
0:  Shapley value of each training data point in $N$
1:  for $k$-th epoch do
2:     Acquire the loss and gradients for each data point in $N$, i.e., $h_{i},\nabla f(i;\theta_{k})$
3:     Using Equation 4 to calculate $\phi_{i}^{k}$, the Shapley value of $i$-th data point in $k$ epoch, where $\alpha=\frac{1}{N}\sum_{i\in N}h_{i}\nabla f(i;\theta)$ and $x_{i}=h_{i}\nabla f(i;\theta_{k})$
4:     Train model on $N$, updating parameters to $\theta_{k+1}$
5:  end for
6:  Return the average Shapley value for each data point across epochs: $\sum_{k=1}^{K}\phi_{i}^{k}/K$
@@ 4 Data Selection
4.1 From Data Valuation to Data Selection
The Shapley value is a widely used method for quantifying individual contributions to overall utility in cooperative games. Previous work (Ghorbani & Zou, 2019; Wang & Jia, 2023) has also heuristically prioritized data points with high Shapley values for model training. In this section, we aim to lay the theoretical groundwork for employing Shapley values in data selection.
Theorem 2
Let $N$ be a set of players and $U$ a utility function. The Shapley value of player $i\in N$ is denoted by $\phi_{i}(U)$. Suppose that for all $i$ and for all subsets $S\subseteq N\setminus\{i\}$, the utility difference satisfies $-m\leq U(S\cup\{i\})-U(S)\leq M$. Then:
  1. 1.
    If $\phi_{i}(U)\geq 0$, then
    $P_{\pi\sim\Pi}\left[U(S_{\pi}^{i}\cup\{i\})\geq U(S_{\pi}^{i})\right]\geq\frac{\phi_{i}(U)}{M},$
  2. 2.
    If $\phi_{i}(U)\leq 0$, then
    $P_{\pi\sim\Pi}\left[U(S_{\pi}^{i}\cup\{i\})\geq U(S_{\pi}^{i})\right]\leq 1+\frac{\phi_{i}(U)}{m}.$
Here, $S_{\pi}^{i}$ represents the set of players preceding player $i$ in a uniform random permutation $\pi$.
The proof can be found in Appendix C.
The theorem offers a probabilistic interpretation of Shapley values. Supposing $S_{\pi}^{i}$ represents the dataset selected so far to some extent, then, if $i$ has a positive Shapley value, the probability that it positively contributes to the coalition $S_{\pi}^{i}$ has a lower bound that increases proportionally with its Shapley value. Conversely, if $i$ has a negative Shapley value, the upper bound on this probability decreases as its Shapley value becomes smaller. Thus, when $i$ has a larger Shapley value, adding it to the selected training subset $S_{\pi}^{i}$ is more likely to yield a positive benefit for $S_{\pi}^{i}$. And when $S_{\pi}^{i}$ can not fairly approximate the dataset selected so far, using the original Shapley value for data selection may become ineffective.
@@ 4.1 From Data Valuation to Data Selection
0:  Training dataset $N$, initial parameters $\theta_{0}$, total epochs $K$, selection interval $R=20$, fraction $a$ of selected data.
0:  Final parameters $\theta_{K}$
1:  Initialize subset $S=\emptyset$.
2:  for $k$-th epoch do
@@
6:           Acquire loss and last-layer gradients for $N_{c}$ (the subset of $N$ with label $c$), i.e., $h_{i},\nabla f(i;\theta_{k})$
7:           Using Equation 4 to calculate Shapley values with $\alpha=\frac{1}{N_{c}}\sum_{i\in N_{c}}h_{i}\nabla f(i;\theta)$ and $x_{i}=h_{i}\nabla f(i;\theta_{k})$
8:           Add top $aN_{c}$ points with highest Shapley values to $S$
@@
10:     end if
11:     Train model on $S$, updating parameters to $\theta_{k+1}$
12:  end for
4.2 CHG Shapley-based Data Selection Algorithm
In this section, we introduce several implementation strategies and practical techniques to improve the scalability and efficiency of CHG Shapley, incorporating common tricks from data selection methods (Killamsetty et al., 2021a; b), and all the compared methods (except Full, Random, and AdaptiveRandom) also adopt these tricks:
Interval Data Selection: Instead of selecting new data at every training step, we perform data selection at intervals. For example, during a 300-epoch training process, data selection is done every 20 epochs, while in the remaining epochs, the previously selected subset is used for training. This approach significantly reduces the computational overhead associated with data selection.
Last-Layer Gradients: Given the high dimensionality of gradients in modern deep learning models, we adopt a last-layer gradient approximation. By focusing solely on the gradients of the last layer, this technique accelerates the computation for all methods.
Per-Class Approximations: To further enhance model accuracy, we apply a per-class approach, running the algorithm separately for each class by considering only the data points from that class in each data selection iteration. This reduces both memory usage and computational cost, improving the overall performance of all methods.
We detail the CHG Shapley Data Selection Algorithm in Algorithm 2, integrating the aforementioned techniques to ensure efficient and scalable data valuation-based data selection.
5 Experiments
We first present the data valuation experiments, followed by the data selection experiments. All experiments were conducted on a server equipped with an Intel(R) Xeon(R) Gold 6326 CPU (2.90 GHz) and an NVIDIA A40 GPU. All methods were implemented using PyTorch.
5.1 Data Valuation Setup and Results
Our CHG Shapley data valuation method (Algorithm 1) was implemented using the OpenDataVal framework (Jiang et al., 2023), and experiments were conducted on the CIFAR-10 embedding dataset. The model used for evaluation was a pre-trained ResNet50 followed by logistic regression, with results reported in Table 1.
The methods compared include Leave-One-Out, Influence Subsample (an efficient approximation of the Influence function (Koh & Liang, 2017)), DVRL (Yoon et al., 2020), Data Banzhaf (Wang & Jia, 2023), AME (Lin et al., 2022), LAVA (Just et al., 2023), and Data OOB (Kwon & Zou, 2023). We do not include Data Shapley and Bata Shapley because these two methods are too slow. We evaluated performance on three tasks: noisy label detection (Table 5), noisy feature detection (Fig. 3), and point removal experiments (Fig. 3). Detailed results comparing these methods can be found in Appendix A.
Across the three tasks, Data OOB and CHG Shapley ranked first and second, respectively. Notably, CHG Shapley completed its evaluation in 18 seconds, whereas Data OOB took over 2 hours. These results demonstrate that CHG Shapley not only achieves competitive data valuation and noise detection performance but also offers significant operational efficiency. This efficiency is driven by the analytic form of the proposed utility function, which evaluates the influence of data subsets on training loss. In the next section, we explore CHG Shapley’s effectiveness in data selection for larger, modern datasets.
@@ 5.1 Data Valuation Setup and Results
Table 2: The accuracy (%) and time (h) comparison. All methods are trained with ResNet-18.
@@
Dataset CIFAR10 CIFAR100 Fraction 0.05 0.1 0.3 0.5 0.7 0.05 0.1 0.3 0.5 0.7 Accuracy (%) Full 95.51 77.56 Random 66.63 77.36 89.93 93.11 93.87 22.58 36.51 62.27 69.70 73.04 AdaptiveRandom 83.26 88.78 93.88 94.58 94.66 48.52 60.58 72.45 74.97 75.70 Glister 85.42 90.52 89.80 90.15 93.36 46.54 63.88 71.88 73.93 75.25 GradMatchPB 84.91 90.14 93.73 94.44 94.54 49.01 62.56 71.27 73.66 74.43 Hardness Shapley 63.07 79.96 93.93 94.63 94.78 37.78 55.21 72.46 75.86 75.77 TracIn (Gradient-Dot) 63.47 81.49 94.13 94.85 94.62 40.21 56.55 72.11 75.01 76.33 CGSV (Gradient-Cosine) 72.97 83.49 92.24 93.68 93.98 44.08 53.87 68.16 74.04 75.59 GradE Shapley (This paper) 83.10 89.80 93.10 93.81 93.58 50.38 60.16 70.53 73.66 74.57 CHG Shapley (This paper) 85.47 91.20 93.73 94.42 94.07 53.20 63.70 72.32 74.59 75.31 Time (h) Full 2.2 2.2 Random 0.1 0.3 0.5 1.8 2.4 0.2 0.5 1.1 0.9 1.6 AdaptiveRandom 0.1 0.4 1.1 1.5 1.8 0.2 0.4 0.9 1.3 2.0 Glister 0.4 0.8 2.2 2.4 2.3 0.4 1.0 3.2 3.2 3.6 GradMatchPB 0.2 0.4 1.0 1.4 2.3 0.3 0.5 1.2 2.2 3.0 Hardness Shapley 0.3 0.6 1.0 1.8 2.4 0.4 0.8 1.3 2.9 3.6 TracIn (Gradient-Dot) 0.3 0.5 1.2 1.8 2.4 0.4 0.6 1.3 2.9 2.4 CGSV (Gradient-Cosine) 0.4 0.5 1.1 1.7 2.4 0.4 0.8 1.2 2.8 2.4 GradE Shapley (This paper) 0.3 0.6 1.2 1.8 2.7 0.4 0.5 1.3 1.9 3.6 CHG Shapley (This paper) 0.3 0.6 1.4 1.7 2.2 0.4 0.5 1.3 2.1 3.5
5.2 Data selection setup
In this experiment, we employed a ResNet-18 model, trained from scratch using stochastic gradient descent with a learning rate of 0.05, momentum of 0.9, weight decay set to 5e-4, and Nesterov acceleration. The model was trained over 300 epochs, with data reselection occurring every 20 epochs. The fraction of selected data varied across 0.05, 0.1, 0.3, 0.5, and 0.7.
In the class imbalance experiments, we set the imbalance ratio to 0.3, meaning that for 30% of the classes, only 10% of their training and validation data was retained. In the noisy label experiments, we set the noise ratio to 0.3, meaning that 30% of the labels in the training set were randomly replaced with labels from the available classes.
The compared methods include Full (training the model with the full dataset), Random (selecting data in the first epoch and never updating the selection), AdaptiveRandom (selecting data randomly in every selection epoch), Glister (Killamsetty et al., 2021b) and GradMatchPB (Killamsetty et al., 2021a) (two coreset selection methods), Hardness Shapley (selecting data with the highest loss), TracIn (Pruthi et al., 2020), and CGSV (Xu et al., 2021) (which use gradient inner product or cosine similarity for data selection as in (Xia et al., 2024)), as well as our proposed GradE Shapley and CHG Shapley. To ensure a fair comparison, the target gradient for TracIn, CGSV, GradE Shapley, and CHG Shapley is the mean gradient of all training data, eliminating the need for an additional validation set. All methods were implemented using the Cords framework (Killamsetty et al., 2021b; a).
Table 3: Comparison of performance under class imbalance setting.
Dataset CIFAR10 CIFAR100 Fraction 0.05 0.1 0.3 0.5 0.7 0.05 0.1 0.3 0.5 0.7 Full 90.37 65.11 AdaptiveRandom 73.64 82.00 88.73 89.61 89.92 36.45 49.44 61.38 63.44 64.27 Glister 67.80 81.47 83.16 82.07 88.63 32.79 48.15 60.86 61.94 62.72 GradMatchPB 75.12 82.92 88.21 89.55 89.71 39.45 51.39 61.54 63.97 64.07 Hardness Shapley 35.57 64.95 87.95 89.13 89.56 25.87 43.34 61.42 63.43 64.31 TracIn (Gradient-Dot) 42.97 68.34 88.45 89.42 89.54 31.81 43.52 61.10 63.54 64.02 CGSV (Gradient-Cosine) 63.09 74.45 87.60 89.72 89.01 32.35 42.58 58.03 63.0 64.38 GradE Shapley (This paper) 67.61 80.98 86.46 87.55 88.55 39.65 48.96 57.81 61.02 62.14 CHG Shapley (This paper) 69.94 83.65 87.69 88.61 88.86 40.31 51.88 60.33 62.90 63.89
Table 4: Comparison of performance under noisy label setting.
Dataset CIFAR10 CIFAR100 Fraction 0.05 0.1 0.3 0.5 0.7 0.05 0.1 0.3 0.5 0.7 Full 79.95 56.67 AdaptiveRandom 57.18 63.21 71.84 75.38 77.41 24.12 32.69 45.36 52.18 53.00 Glister 62.48 71.22 75.51 77.62 77.91 18.75 27.23 46.22 49.38 52.93 GradMatchPB 56.33 69.29 73.50 76.44 76.74 26.40 32.55 43.75 48.47 50.50 Hardness Shapley 21.63 35.41 63.31 75.95 77.28 3.72 7.23 43.16 50.90 51.80 TracIn (Gradient-Dot) 30.72 46.87 70.45 77.14 77.90 28.37 29.70 43.05 50.83 52.29 CGSV (Gradient-Cosine) 70.85 77.48 78.90 78.42 76.84 33.12 42.12 55.31 59.23 57.18 CHG Shapley (This paper) 72.18 83.35 88.57 90.00 90.20 36.75 46.93 60.09 62.59 62.79 GradE Shapley (This paper) 79.35 86.74 90.83 92.03 91.35 44.04 52.14 60.25 63.20 62.92
5.3 Data Selection Results
We present the accuracy and time consumption comparisons in Table 2 under the standard dataset setting. Hardness Shapley also performs competitively at higher fractions (e.g., 0.3), aligning with the findings of Qin et al. (2024). However, when the selection ratio is small (e.g., 0.05), its performance significantly declines, likely due to the difficulty in training the network when all selected data points are particularly hard. TracIn also performs well when the selection ratio is large. CHG Shapley demonstrates particularly strong performance when the selection ratio is small (i.e., 0.05, 0.1), effectively identifying high-value data. Furthermore, its time consumption is comparable to other methods, underscoring the efficiency of CHG Shapley as a data valuation approach.
Table 3 presents the results under the class imbalance dataset setting. Similar to the standard dataset setting, CHG Shapley performs well at smaller selection ratios, and its comparison with GradE Shapley highlights the advantage of incorporating data hardness as part of data value.
The most surprising results are presented in Table 4, under the noisy label dataset setting. GradE Shapley demonstrates a 5%-10% performance improvement over other methods (except CHG Shapley) in this scenario. Even with 10%-30% of the data selected, GradE Shapley can outperform using the full dataset, reinforcing its effectiveness in identifying label noise. In contrast, Hardness Shapley performs notably worse in this setting, especially at smaller selection ratios. In other words, the unreliability of hardness in the presence of label noise leads to CHG Shapley underperforming relative to GradE Shapley.
Overall, the combination of efficiency and accuracy demonstrated across these three experiments establishes CHG Shapley as a practical and powerful tool for data selection and valuation in large-scale datasets with small amounts of label noise. And GradE Shapley excels when handling larger amounts of label noise.
@@ 6 Discussion and Future Work
In the data selection experiments, the introduction of hardness improves the model’s performance when no label noise is present. This suggests that hardness can serve as a proxy for data value to some extent. However, in cases with significant label noise, incorporating hardness into GradE Shapley diminishes its effectiveness. This indicates a need for further investigation into how hardness can be fairly integrated into data value assessments, particularly when it becomes an unreliable metric.
Besides, CHG Shapley performs well at small fractions but struggles with larger ones. This challenge may arise from the difficulties of selecting data based on Shapley values (Wang et al., 2024c), rather than the limitations of CHG Shapley itself. Specifically, once a subset of data $s$ is selected, the Shapley value of the remaining data may change, as the set $S_{\pi}^{i}$ in Equation 2 will include $s$. This affects the effectiveness of the original Shapley values in evaluating the remaining data. If we must recalculate the Shapley values for the remaining data each time a new data point is selected, the time complexity becomes unmanage able even with our proposed method. Thus, we leave the exploration of iterative Shapley value calculations for data selection to future work.